OpenStudy (anonymous):

Am I right? Which of the following is the equation of a line parallel to the line y=4x+1, passing through the point (5,1)? A.) 4x-y=19 B.) 4x+y=-19 C.) -4x-y=19 D.)4x+y=19 Is it D?

OpenStudy (phi):

the line you want has the same slope as the line y= 4x+1 that is in the form y=mx + b m is the slope in this case, m matches up with 4 so the slope is 4

OpenStudy (phi):

your line will go through the point (5,1) y = m x + b but we know m=4, and if x is 5, y is 1 1 = 4*5+b 1=20 + b b= 1-20= -19 so your line is y = 4x -19 rearrange to match one of your choices.